Given that βˆ‘ π‘’π‘˜ ∞ π‘˜=1 converges with π‘’π‘˜ > 0, prove that βˆ‘ βˆšπ‘’π‘˜.π‘’π‘˜+1 ∞ π‘˜=1 also converges. Show that the converse is also true if π‘’π‘˜ is monotonic.


1
Expert's answer
2021-12-20T18:46:10-0500

"\sum \sqrt{u_nu_{n+1}}"


since uk is monotonic, then:

"u_n\/u_{n+1}<1"


"u_nu_{n+1}<u_{n+1}^2"


"\sqrt{u_nu_{n+1}}<u_{n+1}"


so, since "\sum u_{n+1}=\sum u_n+u_{n+1}" converges, series "\sum \sqrt{u_nu_{n+1}}" converges as well